OpenStudy (anonymous):

Find the derivative of s = (4t+3)^3(2t)

OpenStudy (anonymous):

Also the derivative for m(x) = \[2x ^{5}\div(4x ^{2}-5x)^{3}\]

OpenStudy (anonymous):

Just to make sure there is no confusion with the first equation, it is s=\[(4t+3)^{3}(2t)\]

OpenStudy (anonymous):

The first one with m = is another question. The second post is the answer to your question.

OpenStudy (anonymous):

[3(4t+3)^{2}(0)\]?

OpenStudy (anonymous):

\[3(4t+3)^{2}(0)\]?

OpenStudy (anonymous):

That is? You multiply the exponential number with whatever is outside the parentheses which is 1 then you minus one from the exponential number then put the derivative after that which is 4 not 0

OpenStudy (anonymous):

I think I got the answer now. It should be \[24(4t+3)^{2}\]

OpenStudy (anonymous):

\[s=\left( 4t+3 \right)^{3}\left( 2t \right)\] \[\frac{ ds }{ dt }=3\left( 4t+3 \right)^{2}\left( 4 \right)\left( 2t \right)+\left( 4t+3 \right)^{3 }*2\] \[=\left( 4t+3 \right)^{2}[24t+2\left( 4t+3 \right)]=\left( 4t+3 \right)^{2}\left( 32t+6 \right)\]

OpenStudy (anonymous):

if y=uv y'=u'v+uv'